# Breaking the Bottleneck with DiffuApriel: High-Throughput Diffusion LMs with Mamba Backbone

### v2

@@ 1 Introduction
The majority of the state-of-the-art Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023a) are trained with the autoregressive (AR) objective: a Transformer predicts the next token given all previous tokens. While simple and powerful, this paradigm faces two persistent limitations. First, AR decoding is inherently sequential: every generated token depends on the full prefix, and thus inference latency grows linearly with output length. Second, AR Transformers are fundamentally constrained by memory: the attention mechanism incurs quadratic complexity in context length, creating computational and memory bottlenecks during both training and auto-regressive inference (Zhou et al., 2024; Tay et al., 2020). These constraints limit throughput and test-time scaling abilitites of these models, motivating research into alternative architectures.
Diffusion Language Models (DLMs) (Li et al., 2025b) provide a flexible alternative to autoregressive generation. Unlike AR decoders, DLMs iteratively denoise entire corrupted sequences in parallel, enabling non-sequential decoding, partial infilling, and semi-autoregressive generation. This opens the door to faster or more controllable generation, especially when fewer denoising steps suffice for acceptable quality. DLMs hold practical promise for scaling iterative reasoning, whether vertically through multiple sampled traces or horizontally through long, coherent single-trace reasoning, as shown in recent AR models (Radhakrishna et al., 2025a, b; Chen et al., 2025). Because DLMs revisit and refine entire sequences during denoising, they can in principle support self-correction by adjusting intermediate errors using both past and future context.
However, a key bottleneck persists: all current DLMs rely on Transformer backbones, making their iterative denoising at inference highly compute-intensive. Each decoding step redundantly recomputes queries, keys, and values (QKV) for every token and layer, even though key–value (KV) states change only marginally between steps. This exhaustive recomputation introduces severe latency and memory overhead, making inference scale with both sequence length and the number of denoising steps, undermining diffusion’s inherent parallelism. The result is a paradox: diffusion promises flexible text generation, yet its efficiency is constrained by the Transformer backbone. Prior work has aimed to accelerate DLMs via improved objectives (Sahoo et al., 2025; Li et al., 2025a; Liu et al., 2025b; Arriola et al., 2025) or Transformer-compatible KV caching (Wu et al., 2025; Ma et al., 2025; Nguyen-Tri et al., 2025), yet the core architectural limitation remains.
@@
In parallel with advances in diffusion language modeling, state-space models (SSMs) have emerged as a powerful class of sequence mixers with linear-time complexity (Gu et al., 2022; Poli et al., 2023; Fu et al., 2023; Gu & Goel, 2023; Dao & Gu, 2024). Recent studies demonstrate that SSMs and SSM-Transformer hybrids can match or even outperform Transformers on language and multimodal benchmarks while achieving substantially higher inference throughput (Gu & Goel, 2023; Somvanshi et al., 2025; Wang et al., 2025a). Despite their potential, SSMs remain unexplored in DLMs. This raises a key question: can structured recurrence serve as an effective language denoiser while enabling faster inference? Exploring the intersection of efficient SSM backbones and masked discrete diffusion for text is the central motivation of our work.
To this end, we introduce the first family of SSM-based discrete masked diffusion language models. Two observations motivate our approach: diffusion LMs are becoming competitive at scale (Gong et al., 2025; Ye et al., 2025; Labs et al., 2025) yet remain tied to Transformer backbones, and recent advances show that Mamba/SSM architectures enable efficient multi-step diffusion in vision (Hu et al., 2024; Ergasti et al., 2025; Ju & Zhou, 2024; Dang et al., 2024). To our knowledge, no prior work has joined these threads for language. Our study fills this gap by showing that SSMs can serve as drop-in denoisers for masked discrete diffusion, thus substantially improving inference efficiency.
Specifically, we replace the Transformer backbone with (i) fully Mamba-2 (Dao & Gu, 2024) denoisers and (ii) hybrid Mamba-2 + Transformer denoisers that retain full-sequence bidirectionality while improving efficiency. Removing quadratic attention reduces per-step latency without altering the probabilistic semantics of masked diffusion. Moreover, block diffusion methods (Arriola et al., 2025; Wu et al., 2025) are orthogonal to our architectural design and can be naturally integrated: in block-DLM hybrids, local Transformer blocks manage short-range dependencies, while SSM layers propagate information globally, combining local precision with efficient long-range reasoning. We leave this direction for future work.
Our contributions are conceptual and empirical rather than algorithmic:
@@
  • New architectural direction. We propose DiffuApriel, which replaces Transformer denoisers with bidirectional Mamba-2 for discrete masked diffusion language modeling. This demonstrates that iterative denoising does not inherently require attention, establishing linear-time backbones as a compelling alternative for scaling diffusion LMs efficiently.
  • Hybrid design exploration. Beyond using only Mamba-2, we also propose DiffuApriel-H, a hybrid architecture that combines attention and state-space layers. We analyze how localized recurrence and global context interact during the denoising process, while still delivering higher throughput than transformer based DLMs. These results provide new insights into the complementary roles of recurrence and attention in diffusion-based generation.
@@
  • Controlled evaluation across scales. We conduct a systematic comparison between DiffuApriel and Transformer-based diffusion models (DiffuTran) under identical training data, tokenization, noise schedules, and decoding steps at three parameter budgets (240M, 0.5B, and 1.3B). Our experiments demonstrate that DiffuApriel and DiffuApriel-H not only improve modeling quality but also deliver substantially higher inference throughput and lower latency, paving the way for more efficient design choices in DLMs.
@@ 2 Related Work
Diffusion modeling replaces left-to-right decoding with iterative denoising. While early work focused on continuous domains (Ho et al., 2020; Song & Ermon, 2021; Rombach et al., 2022; Peebles & Xie, 2023), recent research has adapted diffusion to discrete text. Austin et al. (2021) introduced structured transition matrices with absorbing states, and DiffusionBERT (He et al., 2023) employed a masked corruption process paired with a BERT-style denoiser. Subsequent studies have further refined masking strategies, noise schedules, and sampling procedures (Chen et al., 2023; Lou et al., 2023; Varma et al., 2025; Fathi et al., 2025). Masked diffusion approaching autoregressive quality (Sahoo et al., 2024, 2025) enabled scalable models (Ye et al., 2025; Nie et al., 2025) competitive with LLaMA (Touvron et al., 2023b), with further gains from AR adaptation and instruction tuning (Gong et al., 2025; Zhu et al., 2025). To enhance inference efficiency, recent acceleration techniques (Ma et al., 2025; Wu et al., 2025; Liu et al., 2025c; Israel et al., 2025; Wang et al., 2025b) use approximate KV-caching mechanisms to speed up DLM inference. In contrast, DiffuApriel-H achieves its speedups without relying on KV caching, though its attention layers remain compatible with these optimizations.
@@
SSMs introduce structured recurrences computable via fast convolutions or scans, achieving linear-time scaling with sequence length. Early models like S4 (Gu et al., 2022) demonstrated strong long-context capacity, while Hyena (Poli et al., 2023) and H3 (Fu et al., 2023) explored alternative structured operators for language modeling. Mamba (Gu & Goel, 2023) extended this line by introducing input-conditioned selective state spaces, matching or surpassing Transformers in accuracy with lower latency and memory use. Somvanshi et al. (2025) further survey the breadth of SSM architectures and highlight their efficiency advantages. In autoregressive language modeling, SSMs rival Transformers at comparable scales while offering higher tokens-per-second throughput. More recent works have evolved various linear-complexity language models and hybrids (Wang et al., 2025a; Ostapenko et al., 2025), ranging from vector recurrences (De et al., 2024) to advanced gating mechanisms (Yang et al., 2024). Yet, these gains remain unexplored in discrete diffusion LMs, where all prior models rely on full-attention denoisers, leaving open whether SSM recurrences can effectively replace attention in iterative denoising for text.
@@ 3.1 Preliminary
$q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})=\prod_{i=1}^{L}q_{t|0}(x_{t}^{i}|x_{0}^{i})=\prod_{i=1}^{L}\text{Cat}(x_{t}^{i};(1-t)\delta_{x_{0}^{i}}+t\delta_{\text{MASK}}),$ (1)
@@
Recent theory (MDLM (Shi et al., 2024; Sahoo et al., 2024), RADD (Ou et al., 2024)) simplifies training from a variational bound to a reweighted cross-entropy over masked positions
$\mathcal{L}_{\text{MDM}}=\int_{0}^{1}\frac{1}{t}\,\mathbb{E}_{q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})}\left[\sum_{i:x_{t}^{i}=\text{MASK}}-\log p_{\theta}(x_{0}^{i}|\mathbf{x}_{t})\right]dt.$ (2)
@@
$q_{s|t}=\prod_{i=0}^{n-1}q_{s|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t}),\quad q_{s|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t})=\begin{cases}1,&\boldsymbol{x}_{t}^{i}\neq[\text{MASK}]\ \text{and}\ \boldsymbol{x}_{s}^{i}=\boldsymbol{x}_{t}^{i}\\ \frac{s}{t},&\boldsymbol{x}_{t}^{i}=[\text{MASK}]\ \text{and}\ \boldsymbol{x}_{s}^{i}=[\text{MASK}]\\ \frac{t-s}{t}\,q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t}),&\boldsymbol{x}_{t}^{i}=[\text{MASK}]\ \text{and}\ \boldsymbol{x}_{s}^{i}\neq[\text{MASK}].\end{cases}$ (3)

### v3

# arXiv:2511.15927v3[cs.LG] 27 Feb 2026

@@ 1 Introduction
The majority of the state-of-the-art Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023a) rely on full multi-head attention mechanism (MHA) (Bahdanau et al., 2014; Vaswani et al., 2017) and are trained with the autoregressive (AR) objective: a Transformer predicts the next token given all previous tokens. While simple and powerful, this paradigm faces several persistent limitations. First, AR decoding is inherently sequential: every generated token depends on the full prefix, and thus inference latency grows linearly with output length. The AR training paradigm also limits data efficiency: each training example provides supervision for only a single next-token prediction, constraining the amount of learning signal that can be extracted from limited data (Ni et al., 2025a). This limitation becomes increasingly relevant as high-quality data, rather than compute, emerges as a key bottleneck in scaling language models. Furthermore, AR Transformers are constrained by both memory and compute: the KV cache grows linearly with context length, inducing increasing memory pressure, while the attention mechanism incurs quadratic compute cost in the sequence length, limiting long-context inference and training (Zhou et al., 2024; Tay et al., 2020). These constraints limit throughput and test-time scaling efficiency, motivating research into alternative architectures.
Diffusion Language Models (DLMs) (Li et al., 2025) provide a flexible alternative to AR generation. Unlike AR decoders, DLMs iteratively denoise entire corrupted sequences in parallel, enabling non-sequential multi-token generation, partial infilling, and self-correction. In principle, this opens the door to faster generation and self-refined reasoning, especially when fewer denoising steps suffice for acceptable quality. However, a key bottleneck persists: all current DLMs rely on Transformer backbones, making their iterative denoising compute or/and memory intensive pushing their throughput well below the AR competitors especially at long sequences. In standard DLMs, denoising proceeds via repeated full sequence re-encoding using bidirectional attention, as each step conditions on both past and future tokens. Because token states evolve across denoising steps, representations cannot be incrementally reused, leading to per-step costs that scale quadratically with sequence length. While several DLM variants enable KV caching (Wu et al., 2025; Ma et al., 2025; Nguyen-Tri et al., 2025; Fathi et al., 2025), the core architectural limitation remains: as the cache grows with sequence length, it induces increasing memory traffic becoming the key bottleneck at inference. Consequently, while inference cost necessarily scales with the number of denoising steps and the sequence length, Transformer-based DLMs incur additional attention and cache related overheads making per token latency grow with sequence length. The result is a paradox: DLMs promise flexible text generation, yet their efficiency is constrained by increasing memory overhead and periodic cache recomputation induced by the Transformer backbone (Wu et al., 2025).
@@
In parallel with advances in diffusion language modeling, state-space models (SSMs) have emerged as a powerful class of sequence mixers with linear-time complexity (Gu et al., 2022; Poli et al., 2023; Fu et al., 2023; Gu & Goel, 2023; Dao & Gu, 2024). Recent studies demonstrate that SSMs and SSM–Transformer hybrids can match or even outperform Transformers while achieving substantially higher inference throughput (Gu & Goel, 2023; Somvanshi et al., 2025; Wang et al., 2025a). Despite their potential, SSMs remain unexplored in DLMs. This raises a key question: can structured recurrence serve as an effective language denoiser while enabling faster inference? Our central motivation is to explore this intersection of efficient SSM backbones and masked discrete diffusion for language modeling.
To this end, we pre-train standard full-sequence DLMs (Sahoo et al., 2024) whose Transformer backbone is modified in two ways: (i) by replacing all Multi-Head Attention (MHA) mixers with Mamba-2 (Dao & Gu, 2024), and (ii) by replacing only a subset of MHA mixers with Mamba-2 to form a hybrid architecture. Further, we employ bidirectional Mamba-2 mixers, as masked diffusion requires conditioning on both past and future context at each denoising step. By removing quadratic attention, we reduce per-step latency and memory pressure without altering the probabilistic semantics of masked diffusion. As a result, Mamba-based DLMs achieve higher throughput than both Transformer-based DLMs and autoregressive baselines across a range of decoding algorithms, including block diffusion (Arriola et al., 2025; Wu et al., 2025).
The main contributions of this work are:
@@
  • New architectural direction. We propose DiffuMamba, which replaces Transformer denoisers with bidirectional Mamba-2 mixers for discrete masked diffusion language modeling. We further introduce DiffuMamba-H, a sparse hybrid variant that interleaves attention by inserting one Transformer block every five Mamba blocks ($\approx 20\%$ attention), starting with an attention block. Together, these models demonstrate that iterative denoising does not inherently require dense attention, positioning linear-time backbones as a scalable alternative for diffusion language models.
@@
  • Controlled evaluation across scales. We conduct a systematic comparison between DiffuMamba and Transformer-based diffusion models (DiffuTran) under identical training data, tokenization, noise schedules, and decoding steps at three parameter budgets (240M, 0.5B, and 1.3B). Our experiments demonstrate that DiffuMamba and DiffuMamba-H can match the modeling quality of the full transformer.
  • Comprehensive throughput benchmarking. We present an extensive asymptotic and empirical analysis of modern DLM inference strategies, scaling generation length beyond 100k tokens, a regime particularly relevant for complex reasoning workloads. Across all evaluated inference algorithms, Mamba-backed DLMs consistently outperform Transformer-based counterparts: by up to 8.2× in full-sequence denoising (LABEL:fig:inf:a and LABEL:fig:inf:b) and by 2.3× in the most efficient block-wise autoregressive denoising setting (LABEL:fig:infb:a and LABEL:fig:infb:b). Notably, the latter is the only inference strategy that scales linearly with sequence length, and when combined with Mamba-based denoisers, it outperforms all other baselines, outlining a promising path for future diffusion-based generation systems.
@@ 2 Related Work
Diffusion modeling replaces left-to-right decoding with iterative denoising. While early work focused on continuous domains (Ho et al., 2020; Song & Ermon, 2021; Rombach et al., 2022; Peebles & Xie, 2023), recent research has adapted diffusion to discrete text. Austin et al. (2021) introduced structured transition matrices with absorbing states, and DiffusionBERT (He et al., 2023) employed a masked corruption process paired with a BERT-style denoiser. Subsequent studies have further refined masking strategies, noise schedules, and sampling procedures (Chen et al., 2023; Lou et al., 2023; Varma et al., 2025; Fathi et al., 2025). Masked diffusion approaching autoregressive quality (Sahoo et al., 2024, 2025) enabled scalable models (Ye et al., 2025; Nie et al., 2025) competitive with LLaMA (Touvron et al., 2023b), with further gains from AR adaptation and instruction tuning (Gong et al., 2025; Zhu et al., 2025). To enhance inference efficiency, recent acceleration techniques (Ma et al., 2025; Wu et al., 2025; Liu et al., 2025b; Israel et al., 2025; Wang et al., 2025b) use approximate KV-caching mechanisms to speed up DLM inference. In contrast, DiffuMamba-H achieves its speedups without relying on KV caching, though its attention layers remain compatible with these optimizations.
@@
SSMs introduce structured recurrences computable via fast convolutions or scans, achieving linear-time scaling with sequence length. Early models like S4 (Gu et al., 2022) demonstrated strong long-context capacity, while Hyena (Poli et al., 2023) and H3 (Fu et al., 2023) explored alternative structured operators for language modeling. Mamba (Gu & Goel, 2023) extended this line by introducing input-conditioned selective state spaces, matching or surpassing Transformers in accuracy with lower latency and memory use. Somvanshi et al. (2025) further survey the breadth of SSM architectures and highlight their efficiency advantages. In autoregressive language modeling, SSMs rival Transformers at comparable scales while offering higher tokens-per-second throughput. More recent works have evolved various linear-complexity LMs and hybrids (Wang et al., 2025a; Ostapenko et al., 2025), ranging from vector recurrences (De et al., 2024) to advanced gating mechanisms (Yang et al., 2024). Although SSM backbones have been used in diffusion models for non-textual domains such as generative modelling in biological sequences (Sahoo et al., 2024), text diffusion LMs continue to rely on full-attention denoisers, leaving open whether SSM recurrences can replace attention for iterative denoising.
@@ 3.1 Preliminary
$q_{t|0}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\prod_{i=1}^{L}q_{t|0}(x_{t}^{i}\mid x_{0}^{i})$
$\prod_{i=1}^{L}\text{Cat}\!\left(x_{t}^{i};\;(1-t)\delta_{x_{0}^{i}}+t\,\delta_{\text{MASK}}\right),$ (1)
@@
Recent theory (MDM (Shi et al., 2024; Sahoo et al., 2024), RADD (Ou et al., 2024)) simplifies training from a variational bound to a reweighted cross-entropy over masked positions
$\small{\mathcal{L}_{\text{MDM}}=\int_{0}^{1}\frac{1}{t}\mathbb{E}_{q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})}\left[\sum_{i:x_{t}^{i}=\text{MASK}}\!\!\!\!\!\!\!\!-\log p_{\theta}(x_{0}^{i}|\mathbf{x}_{t})\right]dt.}$ (2)
@@
$q_{s|t}=\prod_{i=1}^{L}q_{s|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t}).$
$\begin{cases}\qquad\qquad 1&\text{if }\boldsymbol{x}_{s}^{i}=\boldsymbol{x}_{t}^{i}\neq[\text{MASK}],\\[4.0pt] \qquad\qquad\dfrac{s}{t}&\text{if }\boldsymbol{x}_{s}^{i}=\boldsymbol{x}_{t}^{i}=[\text{MASK}],\\[6.0pt] \dfrac{t-s}{t}\,q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t})&\text{if }\boldsymbol{x}_{t}^{i}=[\text{MASK}]\neq\boldsymbol{x}_{s}^{i}.\end{cases}$ (3)